OpenStudy (anonymous):

Find the vertex, the line of symmetry, the maximum of the quadratic function, and the graph function. f(x)=-2x^2+2x+5 What is the x-coordinate of the vertex?

OpenStudy (mertsj):

x = -b/2a x = -2/-4=1/2 y=-2(1/2)^2 +2(1/2) +5 y = -1/2+1+6=5.5 Vertex is (1/2,11/2) Line of symmetry goes through the vertex and is vertical so x = 1/2 The maximum is the y coordinate of the vertex so 11/2 The graph contains the vertex, the y intercept of 5, (2,1) (3,-7). Using symmetry, that should be enough to graph it. The x coordinate of the vertex is 1/2